The value of charge on the oil droplets expt observed were - 1.6 ×10^- 19 and - 4×10^- 19. The value of the electronic charge indicated by these results is?   Ans-          0.8×10^- 19

Grade:12

1 Answers

Vagdevi
5 years ago
The ratio of the charges gives the value of n1&n2.bcz q=ne.
Q1\Q2=1.6\4=2\5.
q\n=e=1.6×10^-19/2=0.8×10^-19